# GNN-based Anchor Embedding for Exact Subgraph Matching

### v1

@@ Abstract.
Subgraph matching query is a classic problem in graph data management and has a variety of real-world applications, such as discovering structures in biological or chemical networks, finding communities in social network analysis, explaining neural networks in machine learning, and so on. To further solve the subgraph matching problem, several recent advanced works attempt to utilize deep-learning-based techniques to handle the subgraph matching query. However, most of these works only obtain approximate results for subgraph matching without theoretical guarantees of accuracy. In this paper, we propose a novel and effective graph neural network (GNN)-based anchor embedding framework (GNN-AE), which allows exact subgraph matching. Unlike GNN-based approximate subgraph matching approaches that only produce inexact results, in this paper, we pioneer the anchor concept (including anchor, anchor graph/path, etc.) in subgraph matching and carefully devise the anchor (graph) embedding technique based on GNN models. We transform the subgraph matching problem into a search problem in the embedding space via the anchor (graph & path) embedding techniques. With the proposed anchor matching mechanism, GNN-AE can guarantee subgraph matching has no false dismissals. We design an efficient matching growth algorithm, which can retrieve the locations of all exact matches in parallel. We also propose a cost-model-based DFS query plan to enhance the parallel matching growth algorithm. Through extensive experiments on 6 real-world and 3 synthetic datasets, we confirm the effectiveness and efficiency of our GNN-AE approach for exact subgraph matching.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2025.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
@@ 1. Introduction
Our Contributions. In this paper, we study the exact subgraph matching problem on the large-scale data graph and propose a novel GNN-AE (GNN-based Anchor Embedding) framework for exact and efficient subgraph matching. In contrast to existing GNN-based approaches for approximate subgraph matching without theoretical guarantees of accuracy, our GNN-AE can obtain exact results for subgraph matching and retrieve the locations of all matches. As shown in Figure 2, in GNN-AE, we first propose the anchor concept (including anchor, anchor graph/path, etc.). We design an effective anchor embedding technique (including anchor graph embedding, and anchor path embedding) based on the anchor concept and GNN models, which transforms the subgraph matching problem into a search problem in the embedding space. If two anchor graphs/paths in the subgraph matching have an isomorphic relationship, they are mapped to the same embedding vector. Based on this principle, we can guarantee a 100% matching (i.e., candidate anchors) recall ratio for each query anchor without introducing false dismissals. GNN-AE does not rely on query graphs to train the model, and the proposed truncated GNN model training technique enables GNN-AE to have a low offline computation cost.
@@ 2.2. GNN-based Subgraph Matching Framework
To address the exact subgraph matching problem, we propose the GNN-based anchor embedding (GNN-AE) framework based on the GNNs technique and the anchor concept (including anchor, anchor graph/path, and anchor (graph/path) embedding, etc.). Algorithm 1 shows the GNN-AE framework, which consists of offline pre-computation and online subgraph matching phases. We pre-process the data graph into anchor (graph/path) embeddings via GNNs and path encoding (lines 1-11), build indexes (lines 12-14) in the offline pre-computation phase, and then answer online subgraph matching queries over indexes (lines 15-21).
@@ 3.4. Matching via Anchor Path Embedding
Figure 13. Illustration of dual 1-hop anchor path encoding.
@@ 29                         Stop the growth of this branch in $S_{t}$
Given an anchor $\psi_{uv}$ in data graph $G$ and an anchor $\psi_{q_{i}q_{j}}$ in query graph $Q$, the anchor $\psi_{uv}$ can be a candidate anchor of $\psi_{q_{i}q_{j}}$, if for any dual 1-hop anchor path $p^{d}_{q_{i}q_{j}}$ of $\psi_{q_{i}q_{j}}$, $\psi_{uv}$ has a dual 1-hop anchor path $p^{d}_{uv}$ aligned with $p^{d}_{q_{i}q_{j}}$.
@@ 7. Related Work
Other Learning-based Graph Computation. In recent years, learning-based methods have emerged in a few related graph computation work, such as subgraph counting (Zhao et al., 2021; Liu et al., 2020; Wang et al., 2022a; Hou et al., 2024), shortest path distance (Huang et al., 2021), graph edit distance computation (Yang and Zou, 2021) and graph keyword search (Hao et al., 2021). However, these graph computation problems (e.g., count or graph edit distance) are different from the subgraph matching problem, and the learning techniques in these works cannot be applied to solve our subgraph matching problem.
@@ 8. Conclusion
In this paper, we propose a novel GNN-based anchor embedding (GNN-AE) framework for exact and efficient subgraph matching. We first propose the anchor concept in subgraph matching. Based on the anchor concept and GNN models, we carefully design effective anchor embedding techniques (including anchor graph embedding, and anchor path embedding), which transforms the subgraph matching problem into a search problem in the embedding space. In addition, a parallel matching growth algorithm is developed to obtain the locations of all exact matches. Extensive experiments have been conducted on 6 real-world and 3 synthetic datasets to show the effectiveness and efficiency of our GNN-AE approach.

### v3

@@ Abstract.
Subgraph matching query is a classic problem in graph data management and has a variety of real-world applications, such as discovering structures in biological or chemical networks, finding communities in social network analysis, explaining neural networks, and so on. To further solve the subgraph matching problem, several recent advanced works attempt to utilize deep-learning-based techniques to handle the subgraph matching query. However, most of these works only obtain approximate results for subgraph matching without theoretical guarantees of accuracy. In this paper, we propose a novel and effective graph neural network (GNN)-based anchor embedding framework (GNN-AE), which allows exact subgraph matching. Unlike GNN-based approximate subgraph matching approaches that only produce inexact results, in this paper, we pioneer a series of concepts related to anchor (including anchor, anchor graph/path, etc.) in subgraph matching and carefully devise the anchor (graph) embedding technique based on GNN models. We transform the subgraph matching problem into a search problem in the embedding space via the anchor (graph & path) embedding techniques. With the proposed anchor matching mechanism, GNN-AE can guarantee subgraph matching has no false dismissals. We design an efficient matching growth algorithm, which can retrieve the locations of all exact matches in parallel. We also propose a cost-model-based DFS query plan to enhance the parallel matching growth algorithm. Through extensive experiments on 6 real-world and 3 synthetic datasets, we confirm the effectiveness and efficiency of our GNN-AE approach for exact subgraph matching.
Artifact Availability:
@@ 1. Introduction
Our Contributions. In this paper, we study the exact subgraph matching problem on the large-scale data graph and propose a novel GNN-AE (GNN-based Anchor Embedding) framework for exact and efficient subgraph matching. In contrast to existing GNN-based approaches for approximate subgraph matching without theoretical guarantees of accuracy, our GNN-AE can obtain exact results for subgraph matching and retrieve the locations of all matches. As shown in Figure 2, in GNN-AE, we first propose a series of concepts related to anchor, including anchor, anchor graph/path, etc. We design effective anchor embedding techniques (including anchor graph embedding, and anchor path embedding) based on anchor concepts and GNN models, which transform the subgraph matching problem into a search problem in the embedding space. If two anchor graphs/paths in the subgraph matching have an isomorphic relationship, they are mapped to the same embedding vector. Based on this principle, we can guarantee a 100% matching (i.e., candidate anchors) recall ratio for each query anchor without introducing false dismissals. GNN-AE does not rely on query graphs to train the model, and the proposed truncated GNN model training technique enables GNN-AE to have a low offline computation cost.
@@ 2.2. GNN-based Subgraph Matching Framework
To address the exact subgraph matching problem, we propose the GNN-based anchor embedding (GNN-AE) framework based on the GNNs technique and the anchor concept (including anchor, anchor graph/path, and anchor (graph/path) embedding, etc). Algorithm 1 shows the GNN-AE framework, which consists of offline pre-computation and online subgraph matching phases. We pre-process the data graph into anchor (graph/path) embeddings via GNNs and path encoding (lines 1-11), build indexes (lines 12-14) in the offline pre-computation phase, and then answer online subgraph matching queries over indexes (lines 15-21).
@@ 3.4. Matching via Anchor Path Embedding
Figure 13. Illustration of dual 1-hop anchor path encoding.
@@ 29                         Stop the growth of this branch in $S_{t}$
Given an anchor $\psi_{uv}$ in data graph $G$ and an anchor $\psi_{q_{i}q_{j}}$ in query graph $Q$, the anchor $\psi_{uv}$ can be a candidate anchor of $\psi_{q_{i}q_{j}}$, if for any dual 1-hop anchor path $p^{d}_{q_{i}q_{j}}$ of $\psi_{q_{i}q_{j}}$, $\psi_{uv}$ has a dual 1-hop anchor path/subpath $p^{d}_{uv}$/$\tau^{d}_{uv}$ aligned with $p^{d}_{q_{i}q_{j}}$.
@@ 8. Conclusion
In this paper, we propose a novel GNN-based anchor embedding (GNN-AE) framework for exact and efficient subgraph matching. We first propose a series of concepts related to anchor (including anchor, anchor graph/path, etc.) in subgraph matching. Based on anchor concepts and GNN models, we carefully design effective anchor embedding techniques (including anchor graph embedding, and anchor path embedding), which transform the subgraph matching problem into a search problem in the embedding space. In addition, a parallel matching growth algorithm is developed to obtain the locations of all exact matches. Extensive experiments have been conducted on 6 real-world and 3 synthetic datasets to show the effectiveness and efficiency of our GNN-AE approach.